a. What is the midpoint of the 2 up to 4 class? (Round the final answer to 1 decimal place.)


b. How many days were 2 up to 4 patients admitted?



c. Approximately how many days were studied?



d. What is the class interval?



e. What is this chart called?
1
Expert's answer
2020-02-26T11:03:42-0500

The following chart shows the number of patients admitted daily to Memorial Hospital

through the emergency room.


a. What is the midpoint of the 2 up to 4 class?

Lower limit "L=2." Upper limit "U=4"


"midpoint=M={L+U \over 2}={2+4 \over 2}=3"


b. How many days were 2 up to 4 patients admitted?

Number of days for which 2 up to 4 patients were admitted is between 20 and 30.


c. Approximately how many days were studied?

30 days were studied.


d. What is the class interval?


"Class\ interval=2-0=2"

e. What is this chart called?

This chart is called frequency polygon.